OpenStudy (anonymous):

Indicate the equation of the given line in standard form. The line containing the hypotenuse of right triangle ABC where A(-5, 5), B(1, 1), and C(3, 4) are the vertices. I worked it out and got x+8y=31 but I'm not sure if its right...

jimthompson5910 (jim_thompson5910):

you're close, but not quite there

jimthompson5910 (jim_thompson5910):

The left side is correct (x+8y), but the right side is NOT correct (it should be some number that is not 31)

OpenStudy (anonymous):

This is how i worked it out....

OpenStudy (anonymous):



OpenStudy (anonymous):

1. midpoint of AB = 1/2(1-5), 1/2(1+5) =-2,3 2. midpoint of BC = 1/2(3+1), 1/2(4+1) = 2,5/2

jimthompson5910 (jim_thompson5910):

why are you looking for midpoints?

OpenStudy (anonymous):

oh my bad. I put the wrong question lol. It's this: Indicate the equation of the given line in standard form. The line containing the midpoints of the legs of right triangle ABC where A(-5, 5), B(1, 1), and C(3, 4) are the vertices.

jimthompson5910 (jim_thompson5910):

ok one sec

jimthompson5910 (jim_thompson5910):

ok you have the correct midpoints

jimthompson5910 (jim_thompson5910):

feel free to keep going

OpenStudy (anonymous):

but do i still have it wrong??

jimthompson5910 (jim_thompson5910):

yes

jimthompson5910 (jim_thompson5910):

the left side of x+8y is correct, but the right side is not

jimthompson5910 (jim_thompson5910):

for now, set the left side equal to some number C x+8y = C Now plug in one midpoint, say (-2, 3) to get x+8y = C -2+8(3) = C ... plug in x = -2 and y = 3 -2 + 24 = C 22 = C C = 22

jimthompson5910 (jim_thompson5910):

So this proves that (-2,3) is on the line x+8y = 22

jimthompson5910 (jim_thompson5910):

you can check the other midpoint, but it should work out as well

jimthompson5910 (jim_thompson5910):

so the line through those two midpoints is x+8y = 22

OpenStudy (anonymous):

Thanks jimmy :)

jimthompson5910 (jim_thompson5910):

you're welcome